The data below are the quiz scores of 10 randomly selected statistics students and the number of hours they studied for the quiz.
Hours,x= 3, 5, 2, 8, 2, 4, 4, 5, 6, 3

Scores, y= 65, 80, 60, 88, 66, 78, 85, 90, 90, 71

Find the correlation coefficient for the given data.

Group of answer choices

0.7165

0.8465

0.0020

5.044
Answer #1

\bar{x} = \frac{\sum x}{n} =\frac{42}{10} = 4.2\\ \bar{y} = \frac{\sum y}{n} =\frac{773}{10} = 77.3

S.D(x) =\sqrt{\frac{1}{n}\sum (x-\bar{x})^2} = \sqrt{\frac{1}{10}\sum (x-4.2)^2} = 1.777639

S.D(y) =\sqrt{\frac{1}{n}\sum (y-\bar{y})^2} = \sqrt{\frac{1}{10}\sum (y-77.3)^2} = 10.59292

covariance(x,y) =\frac{1}{n}\sum (x*y) - \bar{x}\bar{y} = 340 - 4.2*77.3 = 15.94

correlation coefficient (x,y) = \frac{cov(x,y)}{sd(x)*sd(y)} = \frac{15.94}{1.777639*1.059292} =\frac{15.94}{18.83039} = 0.846504

Ans = 0.8465